Is this correct?

1. Multiply (x-2)(x+2); answer: x^2+4?
2. x^4-2x^3+5x^2+10x+12 for value of x=2 ; answer: 52?
3. A chemist has a 25% and a 50% acid solution. How much of each solution should be used to form 200ml of a 35% acid solution? ; answer: 120 ml of 25% and 80 ml of 50%?
I appreciate the help; hopefully I got them right.

Answer by checkley71(8403)

1) (X-2)(X=2)=X^2-4 ANSWER.
2) X^4-2X^3+5X^2+10X+12 WHEN X=2 THEN
2^4-2*2^3+5*2^2+10*2+12
16-2*8+5*4+20+12
16-16+20+20+12=53 ANSWER.
3) .5X+.25(200-X)=.35*200
.5X+50-.25X=70
.25X=70-50
.25X=20
X=20/.25
X=80 ML OF 50% ACID IS REQUIRED.
200-80=120 ML OF 25% SOLUTION IS REQUIRED.
PROOF:
.5*80+.25*120=70
40+30=70
70=70